OpenStudy (anonymous):

I was frustrated at this, my last question and not sure how to do it properly. I just need guidance in any way. I tried grouping but I may have done it incorrectly. Finding the difference of (−4x2 − 7x3 − 2) − (8x3 + 3x − 7x2).

OpenStudy (ivettef365):

first take out the parenthesis and you have -4x^2 -7x^3 -2 -8x^3 -3x + 7x^2 now combine like terms

OpenStudy (anonymous):

this could be incredibly wrong( would not be surprised) -11^5 -2 - 4x

OpenStudy (jdoe0001):

\(\bf (-4x^2-7x^3 - 2) - (8x^3 + 3x - 7x^2)\\ \quad \\ \implies -4x^2 - 7x^3 - 2 - 8x^3 - 3x + 7x^2\\\quad \\\quad\\ -4x^2 + 7x^2- 7x^3 - 8x^3- 2 - 3x \)

OpenStudy (jdoe0001):

keep in mind that \(\bf \large - (8x^3 + 3x - 7x^2) \implies -1 \times (8x^3 + 3x - 7x^2)\)

OpenStudy (ivettef365):

you can only combine terms that are equal, for example you can combine -4x^2 -7x^2 both terms have x^2 so (-4-7= -11) so combined it is -11x^2

OpenStudy (anonymous):

wow I was infact, so wrong on the grouping of like terms.

OpenStudy (anonymous):

so like in the example, the power ^2 had to be also the same to group aswell as the variable x?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

I did not know that key factor! always that the variable ONLY had to be the same hah

OpenStudy (ivettef365):

correct

OpenStudy (jdoe0001):

that's true for multiplication, not so for subtractions or additions

OpenStudy (anonymous):

probably where my confusing made its way in.

OpenStudy (anonymous):

so, the combined like terms come together as -11x^2? @ivettef365 I got the steps from @jdoe0001 and need to make sure im not lost here.

OpenStudy (jdoe0001):

\(\bf \large \color{red}{-4x^2 + 7x^2}\color{blue}{- 7x^3 - 8x^3}- 2 - 3x\)

OpenStudy (ivettef365):

sorry it is -4x^2 +7x^2 so it is (-4+7= 3) so it is 3x^2

OpenStudy (ivettef365):

now combine the numbers with x^3

OpenStudy (anonymous):

@jdoe0001 do I combine here aswell?

OpenStudy (jdoe0001):

you combine, the like-terms, the terms whose variables are equal in letter and exponent

OpenStudy (jdoe0001):

for example, -2 has no other constant equal in the surroundings, so it's left by its lonesome -3x has no other term in with an "x" only, so it's also left by its lonesome the other 2 pair, have companions

OpenStudy (anonymous):

3x2 - 15x3 -2-3x so.. uhm where do i go wrong.

OpenStudy (jdoe0001):

looks good

OpenStudy (anonymous):

so thats at its simplest form?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

wooooot!

OpenStudy (jdoe0001):

heheh, anything wrong with it? heh

OpenStudy (anonymous):

Nope I think it makes sense :DD!!! THANK YOUU, praise the lorddd

OpenStudy (jdoe0001):

yw